OpenStudy (jpsmarinho):

How to proof this? Is this some property? (L + h)^3 = L^3*(1+ (h/L) )^3

OpenStudy (anonymous):

Yes. Do you know the property of exponents:\[(ab)^n=\left ( a^n\right ) \left ( b^n\right ) \]???

OpenStudy (anonymous):

First,\[h=(h)(1)=(h)\left ( \frac{L}{L}\right ) =(L)\left ( \frac{h}{L}\right )\]So,\[L+h=L+L\left ( \frac{h}{L}\right ) =(L)\left ( 1+\frac{h}{L}\right )\]

OpenStudy (anonymous):

Then apply the property.